OpenStudy (anonymous):

a bag contains 57 coins, with a total weight of 396 grams. if the bag contains only gold coins that weigh 8 grams each and silver coins that weigh 5 grams each, how many gold and how many silver coins are in the bag?

OpenStudy (amistre64):

g+s = 57 8g+5s=396

OpenStudy (amistre64):

(8-5)g = 57(-5)+396